Part of the reason that conventional explosives cause so much damage is that their detonation produces a strong shock wave that can knock things down. While using explosives to knock down a building, the shock wave can be so strong that 12 liters of gas will reach a pressure of 3.8 x 104 mmHg. When the shock wave passes and the gas returns to a pressure of 760 mm Hg, what will the volume of that gas be?


Law: ______________________


Answer:____________________



A hydrogen gas volume thermometer has a volume of 100.0 cm3 when immersed in an ice-water bath at 0°C. When immersed in boiling liquid chlorine, the volume of the hydrogen at the same pressure is 87.2 cm3. Find the temperature of the boiling point of chlorine in °C. (Remember cm3=mL=cc)


Law: ______________________


Answer:____________________


1
Expert's answer
2021-04-28T06:17:33-0400

Boyle's Law


"P_1V_1=P_2V_2"


"12\u00d7[3.8\u00d710^4]=760V_2"


"V_2=600litres"




Charles Law


"\frac{V_1}{T_1} = \frac{V_2}{T_2}"


"\frac{100}{273}=\frac{87.2}{T_2}"


"T_2 =\frac{273×87.2}{100}=238.06 K" or. "-34.94\u00b0c"